# On the Privacy, Security, and Trustworthy for Distributed Wireless Large AI Model (WLAM)

### v1

@@ Abstract
Combining wireless communication with large artificial intelligence (AI) models can open up a myriad of novel application scenarios. In sixth generation (6G) networks, ubiquitous communication and computing resources allow large AI models to serve democratic large AI models-related services to enable real-time applications like autonomous vehicles, smart cities, and Internet of Things (IoT) ecosystems. However, the security considerations and sustainable communication resources limit the deployment of large AI models over distributed wireless networks. This paper provides a comprehensive overview of privacy, security, and trustworthy for distributed wireless large AI model (WLAM). In particular, the detailed privacy and security are analysis for distributed WLAM is fist revealed. The classifications and theoretical findings about privacy and security in distributed WLAM are discussed. Then the trustworthy and ethics for implementing distributed WLAM are described. Finally, the comprehensive applications of distributed WLAM is provided in the aspect of electromagnetic signal processing.
@@ I-A Motivation
The integration of wireless communication and artificial intelligence (AI) technologies has revolutionized numerous sectors [1], from healthcare [2] to transportation [3], by facilitating the development of distributed wireless large AI models (WLAM) [4, 5]. The WLAM includes two aspects, large AI models over wireless networks, and large AI models for solving wireless communication problems. The large AI models leverage the distributed nature of wireless networks to harness computational resources across diverse devices, enabling sophisticated data processing and decision-making in real time [6]. However, the proliferation of WLAM presents unique challenges, particularly regarding privacy, security, and trustworthiness.
@@ I-B Classifications of WLAM
  • Data Distribution: This approach involves distributing datasets across multiple devices. Federated learning is a prime example, allowing devices to independently train local models and share updates with a central server [11]. By transmitting only model updates rather than raw data, this method enhances privacy [12]. However, it relies on sophisticated aggregation algorithms to ensure accurate global model convergence without leaking sensitive information. This balance between privacy and performance is a critical focus of ongoing research [13].
  • Model Distribution: In contrast, model distribution involves dividing the large AI model itself across different devices. Split learning exemplifies this approach, where segments of the model are trained on separate devices [14]. This method splits the original whole model into multiple small models and reduces computational load on individual devices. However, model distribution method requires secure protocols for transmitting intermediate data and integrating model components, as partial data exchanges can expose vulnerabilities [15]. Ensuring the seamless integration of distributed model components remains a significant technical challenge.
@@ II-A Privacy Analysis
In WLAM, the model pret-raining and fine-tuning require data collection and processing. Although the original data can be stored locally, the training data transmitted over wireless communication in WLAM systems may still be vulnerable to interception, allowing an attacker to potentially recover the original data from these transmissions and jeopardize privacy. For example, in federated learning, local model parameters are uploaded from the client to the server [16], and in split learning, partially processed data is transmitted from the client network to the main server-side network, both of which are vulnerable to privacy risks.
@@
  • Encryption Algorithm: To address privacy issues in WLAM, encryption techniques like differential privacy, multi-party computation, homomorphic encryption, and functional encryption are used. In [20], a secure federated learning approach is implemented by combining fully homomorphic encryption with FL in IoT-enabled smart cities, achieving strong securityalong with high accuracy, low communication overhead, and reduced latency.
@@ IV Applications with EM Signal Processing
Electromagnetic (EM) signal processing [50, 51, 52] can significantly enhance the privacy, security, and trustworthiness of distributed WLAM. By leveraging EM properties [53], WLAM systems can achieve secure communication, robust privacy measures, and reliable system integrity, as illustrated in Fig. 4. This section delves into these aspects, highlighting their impact and potential advancements.
@@
The future of WLAM systems will increasingly depend on advancements in EM signal processing to address evolving privacy and security challenges. Emerging technologies like terahertz communication and quantum cryptography offer promising avenues for further enhancing security protocols. These innovations have the potential to improve the speed, accuracy, and robustness of WLAM applications, ensuring that they remain secure and trustworthy as they adapt to new technological landscapes.

### v2

# On Privacy, Security, and Trustworthiness in Distributed Wireless Large AI Models (WLAM)

@@ Abstract
Combining wireless communication with large artificial intelligence (AI) models can open up a myriad of novel application scenarios. In sixth generation (6G) networks, ubiquitous communication and computing resources allow large AI models to serve democratic large AI models-related services to enable real-time applications like autonomous vehicles, smart cities, and Internet of Things (IoT) ecosystems. However, the security considerations and sustainable communication resources limit the deployment of large AI models over distributed wireless networks. This paper provides a comprehensive overview of privacy, security, and trustworthy for distributed wireless large AI model (WLAM). In particular, a detailed privacy and security are analysis for distributed WLAM is fist revealed. The classifications and theoretical findings about privacy and security in distributed WLAM are discussed. Then the trustworthy and ethics for implementing distributed WLAM are described. Finally, the comprehensive applications of distributed WLAM are presented in the context of electromagnetic signal processing.
@@ I-A Motivation
The integration of wireless communication and artificial intelligence (AI) technologies has revolutionized numerous sectors [1], from healthcare [2] to transportation [3], by facilitating the development of distributed wireless large AI models (WLAM) [4, 5]. WLAM encompasses two aspects: large AI models deployed over wireless networks and large AI models designed to solve wireless communication problems. Large AI models leverage the distributed nature of wireless networks to harness computational resources across diverse devices, enabling advanced data processing and real-time decision-making [6]. However, the proliferation of WLAM introduces unique challenges, particularly in the areas of privacy, security, and trustworthiness.
@@ I-B Classifications of WLAM
  • Data Distribution: This approach involves distributing datasets across multiple devices. Federated learning is a prime example, enabling devices to independently train local models and share updates with a central server [11]. By transmitting only model updates rather than raw data, this method enhances privacy [12]. However, it relies on sophisticated aggregation algorithms to ensure accurate global model convergence without leaking sensitive information. This balance between privacy and performance is a critical focus of ongoing research [13].
  • Model Distribution: In contrast, model distribution involves dividing the large AI model itself across different devices. Split learning exemplifies this approach, where segments of the model are trained on separate devices [14]. This method divides the original model into smaller sub-models, reducing the computational load on individual devices. However, model distribution method requires secure protocols for transmitting intermediate data and integrating model components, as partial data exchanges can expose vulnerabilities [15]. Ensuring the seamless integration of distributed model components remains a significant technical challenge.
@@ II-A Privacy Analysis
In WLAM, model pre-training and fine-tuning require extensive data collection and processing. Although original data can be stored locally, the training data transmitted over wireless communication in WLAM systems remains vulnerable to interception, potentially allowing attackers to recover the original data and compromise privacy. The distributed WLAM system involves significant data interactions between devices and servers, increasing the risk of privacy breaches. For example, in federated learning, local model parameters are uploaded from the client to the server [16], and in split learning, partially processed data is transmitted from the client network to the main server-side network, both of which are vulnerable to privacy risks.
@@
  • Encryption Algorithm: To address privacy issues in WLAM, encryption techniques like differential privacy, multi-party computation, homomorphic encryption, and functional encryption are used. Differential privacy methods can be employed to ensure that individual data cannot be inferred from the aggregate outputs of the model. Homomorphic encryption, on the other hand, enables computation on encrypted data, eliminating the need to share raw data with servers.
    In [20], a secure federated learning approach is implemented by combining fully homomorphic encryption with FL in IoT-enabled smart cities, achieving strong securityalong with high accuracy, low communication overhead, and reduced latency.
@@ II-B1 Attack Types and Challenges for WLAM
Data Poisoning Attacks
For example, data poisoning attacks, in which malicious actors inject false data to mislead the AI model, pose a significant threat to system reliability. A key countermeasure against such attacks involves implementing robust data validation mechanisms to identify and isolate malicious contributions in WLAM.
@@ IV Applications with EM Signal Processing
The distributed nature of WLAM offers significant flexibility and robustness for electromagnetic (EM) signal processing. For instance, WLAM can enable real-time beamforming by leveraging the computational power of edge devices to optimize wireless signal transmissions. Moreover, WLAM supports dynamic spectrum management in wireless networks by predicting spectrum availability and efficiently allocating resources.
Besides, EM signal processing [50, 51, 52] can significantly enhance the privacy, security, and trustworthiness of distributed WLAM. By leveraging EM properties [53], WLAM systems can achieve secure communication, robust privacy measures, and reliable system integrity, as illustrated in Fig. 4. This section delves into these aspects, highlighting their impact and potential advancements.
@@
The future of WLAM systems will increasingly depend on advancements in EM signal processing to address evolving privacy and security challenges. Emerging technologies like terahertz communication and quantum cryptography offer promising avenues for further enhancing security protocols. These innovations have the potential to improve the speed, accuracy, and robustness of WLAM applications, ensuring that they remain secure and trustworthy as they adapt to new technological landscapes. Further research is required to develop energy-efficient algorithms capable of operating seamlessly across diverse devices in distributed WLAM systems.
@@ V Conclusions and Future Directions
In conclusion, WLAM offers a promising framework for integrating AI with wireless communication systems, though it necessitates addressing inherent security and privacy challenges. Advancements in cryptographic techniques and distributed algorithms will be pivotal in overcoming these challenges and unlocking the full potential of WLAM. Despite its potential, significant challenges remain for the widespread deployment of distributed WLAM. Scalability remains a challenge, particularly as the number of participating devices increases. Besides, ensuring low-latency communication among distributed components while maintaining data security is a critical challenge.